OpenStudy (anonymous):

.please help! will give medal!! a plane flies 420 miles with the wind and 350 against the wind in the same length of time. if the speed of the wind is 15mph, find the speed of the plane in still air.

OpenStudy (mathstudent55):

v = d/t d = vt the plane speed with the wind is v + 15; the distance is 420 miles the plane speed against the wind is v - 15; the distance is 350 miles 420 = (v + 15)t 350 = (v - 15)t Solve the equations simultaneously and find v.

OpenStudy (jdoe0001):

r = the plane's rate at 420 the wind rate gets added to it at 350 the wind rate gets substracted 420 = r+15 350 = r-15 so, just solve by substitution